STUDIES ON THE PHOTO-DISSOCIATION AND SYMMETRY OF THE D̃ STATE OF SO_{2}^{+}

\underline{ZHONG WANG}^{a}, LIMIN ZHANG, SHILIN LIU, SHUQIN YU, AND XINGXIAO MA, Laboratory of Bond Selective Chemistry, Department of Chemical Physics, University of Science and Technology of China,People’s Republic of China.

Pure parent ions, SO_{2}^{+} X̃^{2}A_{1}(000), were prepared by [3+1] multiphoton ionization of SO_{2} molecules at 380.85 nm. The photodissociation process and the symmetry of the excited states of SO_{2}^{+} molecular ions was investigated by measuring the photofragment SO^{+} excitation (PHOFEX) spectrum in ultra-violet (UV) (282-332 nm) and the visible (562-660 nm) wavelength ranges, respectively. The D̃(v_{1}00), D̃(0v_{2}0) ← X̃^{2}A_{1}(000) transitions of SO_{2}^{+} were assigned in the PHOFEX spectrum in the UV range. By comparing the resolved PHOFEX spectrum in UV range to the continuous PHOFEX spectrum in visible range, it is deduced that around the D̃ state of SO_{2}^{+} there exists a repulsive state (α^{2}A_{2}) converging to the dissociation limit of SO^{+}(X^{2}π) + O(^{3}P_{g}), and the coupling between SO_{2}^{+} D̃ and SO_{2}^{+}(α^{2}A_{2}) leads to the dissociation to SO^{+}(X^{2}π) + O(^{3}P_{g}). The symmetry of the D̃ state is D̃^{2}B_{1}.

aPresent Address: Department of Chemistry, Brookhaven National Laboratory, Upton, NY 11973-5000